Jack Hammer invests in a stock that will pay dividends of $3.09 at the end of the first year; $3.48 at the end of the second year; and $3.87 at the end of the third year. Also, he believes that at the end of the third year he will be able to sell the stock for $59.

What is the present value of all future benefits if a discount rate of 9 percent is applied? Use Appendix B for an approximate answer, but calculate your final answer using the formula and financial calculator methods.

Dividend Present Value
$3.09 ?
3.48 ?
3.87 ?
59.00 ?
Total
Answer #1

Dividend

Present Value

$3.09

$2.83

$3.48

$2.93

$3.87

$2.99

$59.00

$45.56

TOTAL

$54.31

Present Value of Dividend in Year 1

= D1/(1 + r)1

= $3.09/(1 + 0.09)1

= $3.09 / 1.09

= $2.83

Present Value of Dividend in Year 2

= D2/(1 + r)2

= $3.48/(1 + 0.09)2

= $3.48 / 1.18810

= $2.93

Present Value of Dividend in Year 3

= D3/(1 + r)3

= $3.87/(1 + 0.09)3

= $3.87 / 1.29503

= $2.99

Present Value of Stock Price in Year 3

= P3(1 + r)3

= $59.00/(1 + 0.09)3

= $59.00 / 1.29503

= $45.56